# Horseshoe Forests for High-Dimensional Causal Survival Analysis

### v1

@@ Abstract
We develop a Bayesian tree ensemble model to estimate heterogeneous treatment effects in censored survival data with high-dimensional covariates. Instead of imposing sparsity through the tree structure, we place a horseshoe prior directly on the step heights to achieve adaptive global-local shrinkage. This strategy allows flexible regularisation and reduces noise. We develop a reversible jump Gibbs sampler to accommodate the non-conjugate horseshoe prior within the tree ensemble framework. We show through extensive simulations that the method accurately estimates treatment effects in high-dimensional covariate spaces, at various sparsity levels, and under non-linear treatment effect functions. We further illustrate the practical utility of the proposed approach by a re-analysis of pancreatic ductal adenocarcinoma (PDAC) survival data from The Cancer Genome Atlas.
@@ 1 Introduction
Global-local shrinkage priors provide a fully Bayesian and computationally attractive alternative to exact sparsity-inducing methods (e.g. Mitchell and Beauchamp (1988) and Liu et al. (2021)). Originally developed for sparse regression problems (Carvalho et al., 2009), these priors have gained popularity in various contexts, including dynamic shrinkage models (Kowal et al., 2019), graphical models (Li et al., 2019b), and Bayesian neural network compression (Louizos et al., 2017). The class of global-local shrinkage priors includes a wide variety of formulations (Tipping, 2001, Johnstone and Silverman, 2004, Griffin and Brown, 2005, Park and Casella, 2008, Bhadra et al., 2019a). Among these, the horseshoe prior (Carvalho et al., 2009) has emerged as a particularly attractive choice because of its strong theoretical properties and empirical performance in high-dimensional settings (Datta and Ghosh, 2013, van der Pas et al., 2014, 2017a, 2017b, Bhadra et al., 2019b, Hahn et al., 2020). We focus on the horseshoe prior to introduce continuous shrinkage directly on the tree step heights. This approach retains all covariates while adaptively shrinking their contributions. The proposed methodology is readily extendable to other global-local shrinkage priors that allow Gibbs sampling and could be further generalised to incorporate different shrinkage strategies with additional computational effort.
@@ 3 Bayesian regression trees
In this section, we describe how we adapt the Bayesian Additive Regression Trees (BART) prior (Chipman et al., 2010) to high-dimensional covariates by introducing a horseshoe prior on the step heights. This approach shifts regularisation from the tree structure (e.g. Linero (2018)) to the step height parameters. We start with an overview of how BART is applied to model log survival times under the accelerated failure time (AFT) framework. We detail our novel adaptation using a global-local shrinkage prior. We then explain how the model accommodates censored outcomes. Finally, we discuss default hyperparameter choices.
@@ 3.2 Horseshoe prior on the step heights
We present our novel formulation of step height prior based on a global-local shrinkage prior. Given a tree $\mathcal{T}$ with $L$ leaf nodes, we specify a prior on the set of step heights =$\mathcal{H}=\{h_{1},h_{2},\dots,h_{L}\}$. To achieve shrinkage in high-dimensional covariate spaces, we use a class of continuous shrinkage priors on the set of step heights. This class can be represented as a scale mixture of normal distributions. In general form, the prior $p_{\mathcal{H}}$ is specified hierarchically:
@@
This formulation falls within the class of global-local shrinkage priors (Polson and Scott, 2011). The global shrinkage parameter $\tau$ governs the overall degree of shrinkage, encouraging all step heights to be small, while the local shrinkage parameters $\lambda_{\ell}$ allow individual step heights $h_{\ell}$ to deviate from this global tendency when supported by the data. In the tree-based context, the local shrinkage parameters refer to the values assigned to individual leaf nodes, each associated with a distinct region of the covariate space, whereas the global shrinkage parameter is shared across all leaf nodes. This hierarchical structure enables the model to suppress noise in regions lacking signal while preserving flexibility in regions where effects are present. The horseshoe prior is particularly well-suited for this task, as it can adapt to varying degrees of sparsity without imposing excessive shrinkage on substantial signals (van der Pas et al., 2017a, b, Bhadra et al., 2019b).
@@ 4 Posterior inference
    Reversible jump step. We detail the reversible jump Metropolis–Hastings step for updating a single tree. Our approach introduces a novel proposal mechanism for the step heights and local shrinkage parameters. These candidate values are drawn from pseudo-Gibbs updates informed by the parent node, which improves efficiency and mixing.
@@ 4.2 Reversible jump step for updating a tree
The tree structure is jointly updated together with its corresponding set of step heights. We propose a new pair $(\mathcal{T}_{j},\mathcal{H}_{j})$, which is then accepted or rejected using a Metropolis–-Hastings step with a reversible jump move (Green, 1995). The conditional posterior distribution of a single tree factorises as
@@
$\xi$ $\sim\mathcal{IG}\left(\frac{1}{2},\frac{1}{\alpha^{2}}\right)$ $\nu_{\ell}$ $\sim\mathcal{IG}\left(\frac{1}{2},\frac{1}{\alpha^{2}}\right).$
@@ 4.3 Full conditional update of step height parameters
Since $\mathcal{D}^{\top}\mathcal{D}$ is a diagonal matrix with leaf sizes on its diagonal, let $n_{\ell}$ denote the number of observations in leaf $\ell$ and $\bar{\mathcal{R}}_{\ell}$ the sum of residuals in leaf $\ell$. We can then express the conditional posterior of each $h_{\ell}$ as
@@ 7 Discussion
This work introduces a novel regularisation strategy for Bayesian regression tree ensembles by applying shrinkage directly to the step heights via the horseshoe prior. We use continuous global-local shrinkage to adaptively control model complexity in high-dimensional settings, while retaining all covariates in the model. This provides an alternative to methods that exclude covariates from the model and risk omitting important confounders. Our method captures complex non-linear effects and higher-order interactions in both the prognostic and treatment effect functions. These are essential for modelling heterogeneous treatment responses in survival data.
@@ Acknowledgement
where $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ and
@@
The results are presented in Figures 12(a) and 12(b). Under both the Gumbel and logistic error distributions, all methods experience reduced coverage levels. Nevertheless, the Causal Horseshoe Forest consistently achieves higher coverage for the CATE and for the ATE in non-linear scenarios, remaining close to nominal levels even as the dimensionality increases to $p=5000$. The Causal Horseshoe Forest maintains CATE coverage rates around 80% in non-linear settings, while the AFT-BART models exhibit substantial drops, with coverage often falling below 60% or even lower at higher dimensions. The non-parametric AFT-BART version demonstrates slightly greater robustness than the semiparametric variant but still struggles to maintain adequate coverage, especially for heterogeneous treatment effects in high-dimensional scenarios. These findings align with the patterns observed in the main simulations, underscoring the stability and flexibility of the Causal Horseshoe Forest under different error structures and challenging high-dimensional conditions.
@@
The results are presented in Figures 12(a) and 12(b). All methods experience reduced coverage levels under both the Gumbel and logistic error distributions Nevertheless, the Causal Horseshoe Forest consistently achieves higher coverage for the CATE and for the ATE in non-linear scenarios, remaining close to nominal levels even as the dimensionality increases to $p=5000$. In particular, the Causal Horseshoe Forest maintains CATE coverage rates around 80% in non-linear settings, while the AFT-BART models exhibit substantial drops, with coverage often falling below 40% or even lower at higher dimensions. The non-parametric AFT-BART version demonstrates slightly greater robustness than the semiparametric variant but still struggles to maintain adequate coverage, especially for heterogeneous treatment effects in high-dimensional scenarios. These findings align with the patterns observed in the main simulations, underscoring the stability and flexibility of the Causal Horseshoe Forest in challenging high-dimensional settings.

### v2

@@ Abstract
We develop a Bayesian tree ensemble model to estimate heterogeneous treatment effects in censored survival data with high-dimensional covariates. Instead of imposing sparsity through the tree structure, we place a horseshoe prior directly on the step heights to achieve adaptive global–local shrinkage. This strategy allows flexible regularisation and reduces noise. We develop a reversible jump Gibbs sampler to accommodate the non-conjugate horseshoe prior within the tree ensemble framework. We show through extensive simulations that the method accurately estimates treatment effects in high-dimensional covariate spaces, at various sparsity levels, and under non-linear treatment effect functions. We further illustrate the practical utility of the proposed approach by a re-analysis of pancreatic ductal adenocarcinoma (PDAC) survival data from The Cancer Genome Atlas.
@@ 1 Introduction
Global–local shrinkage priors provide a fully Bayesian and computationally attractive alternative to exact sparsity-inducing methods (e.g. Mitchell and Beauchamp (1988) and Liu et al. (2021)). Originally developed for sparse regression problems (Carvalho et al., 2009), these priors have gained popularity in various contexts, including dynamic shrinkage models (Kowal et al., 2019), graphical models (Li et al., 2019b), and Bayesian neural network compression (Louizos et al., 2017). The class of global–local shrinkage priors includes a wide variety of formulations (Tipping, 2001, Johnstone and Silverman, 2004, Griffin and Brown, 2005, Park and Casella, 2008, Bhadra et al., 2019a). Among these, the horseshoe prior (Carvalho et al., 2009) has emerged as a particularly attractive choice because of its strong theoretical properties and empirical performance in high-dimensional settings (Datta and Ghosh, 2013, van der Pas et al., 2014, 2017a, 2017b, Bhadra et al., 2019b, Hahn et al., 2020). We focus on the horseshoe prior to introduce continuous shrinkage directly on the tree step heights. This approach retains all covariates while adaptively shrinking their contributions. The proposed methodology is readily extendable to other global–local shrinkage priors that allow Gibbs sampling and could be further generalised to incorporate different shrinkage strategies with additional computational effort.
@@ 3 Bayesian regression trees
In this section, we describe how we adapt the Bayesian Additive Regression Trees (BART) prior (Chipman et al., 2010) to high-dimensional covariates by introducing a horseshoe prior on the step heights. This approach shifts regularisation from the tree structure (e.g. Linero (2018)) to the step height parameters. We start with an overview of how BART is applied to model log survival times under the accelerated failure time (AFT) framework. We detail our novel adaptation using a global–local shrinkage prior. We then explain how the model accommodates censored outcomes. Finally, we discuss default hyperparameter choices.
@@ 3.2 Horseshoe prior on the step heights
We present our novel formulation of step height prior based on a global–local shrinkage prior. Given a tree $\mathcal{T}$ with $L$ leaf nodes, we specify a prior on the set of step heights =$\mathcal{H}=\{h_{1},h_{2},\dots,h_{L}\}$. To achieve shrinkage in high-dimensional covariate spaces, we use a class of continuous shrinkage priors on the set of step heights. This class can be represented as a scale mixture of normal distributions. In general form, the prior $p_{\mathcal{H}}$ is specified hierarchically:
@@
This formulation falls within the class of global–local shrinkage priors (Polson and Scott, 2011). The global shrinkage parameter $\tau$ governs the overall degree of shrinkage, encouraging all step heights to be small, while the local shrinkage parameters $\lambda_{\ell}$ allow individual step heights $h_{\ell}$ to deviate from this global tendency when supported by the data. In the tree-based context, the local shrinkage parameters refer to the values assigned to individual leaf nodes, each associated with a distinct region of the covariate space, whereas the global shrinkage parameter is shared across all leaf nodes. This hierarchical structure enables the model to suppress noise in regions lacking signal while preserving flexibility in regions where effects are present. The horseshoe prior is particularly well-suited for this task, as it can adapt to varying degrees of sparsity without imposing excessive shrinkage on substantial signals (van der Pas et al., 2017a, b, Bhadra et al., 2019b).
@@ 4 Posterior inference
    Reversible jump step. We detail the reversible jump Metropolis–Hastings step for updating a single tree. Our approach introduces a novel proposal mechanism for the step heights and local shrinkage parameters. These candidate values are drawn from pseudo Gibbs updates informed by the parent node, which improves efficiency and mixing.
@@ 4.2 Reversible jump step for updating a tree
The tree structure is jointly updated together with its corresponding set of step heights. We propose a new pair $(\mathcal{T}_{j},\mathcal{H}_{j})$, which is then accepted or rejected using a Metropolis–-Hastings step with a reversible jump move (Green, 1995). The conditional posterior distribution of a single tree factorises as:
@@
$\xi$ $\sim\mathcal{IG}\left(\frac{1}{2},\frac{1}{\alpha^{2}}\right)$ $\nu_{\ell}$ $\sim\mathcal{IG}\left(\frac{1}{2},\frac{1}{\alpha^{2}}\right),$
@@ 4.3 Full conditional update of step height parameters
Since $\mathcal{D}^{\top}\mathcal{D}$ is a diagonal matrix with leaf sizes on its diagonal, let $n_{\ell}$ denote the number of observations in leaf $\ell$ and $\bar{\mathcal{R}}_{\ell}$ the sum of residuals in leaf $\ell$. We can then express the conditional posterior of each $h_{\ell}$ as:
@@ 7 Discussion
This work introduces a novel regularisation strategy for Bayesian regression tree ensembles by applying shrinkage directly to the step heights via the horseshoe prior. We use continuous global–local shrinkage to adaptively control model complexity in high-dimensional settings, while retaining all covariates in the model. This provides an alternative to methods that exclude covariates from the model and risk omitting important confounders. Our method captures complex non-linear effects and higher-order interactions in both the prognostic and treatment effect functions. These are essential for modelling heterogeneous treatment responses in survival data.
@@ Acknowledgement
The software implementation is available on CRAN: https://cran.r-project.org/package=ShrinkageTrees. All data and replication code can be found at: https://github.com/tijn-jacobs/ShrinkageTrees.
Funding
@@
where $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ and:
@@
The results are presented in Figures 12(a) and 12(b). Coverage drops for all methods under both the Gumbel and logistic error distributions. The Causal Horseshoe Forest consistently performs better. It maintains higher coverage for both the CATE and ATE, especially in non-linear settings. Even at $p=5000$, coverage stays close to the nominal level. Causal Horseshoe Forest keeps CATE coverage around 80% in non-linear scenarios. In contrast, AFT-BART shows a sharp decline. Its coverage often falls below 60%, especially in high dimensions. The non-parametric AFT-BART is slightly more robust than its semiparametric counterpart. Still, it fails to maintain reliable coverage, particularly when treatment effects vary. These results mirror the patterns seen in the main simulations. They highlight the stability and adaptability of the Causal Horseshoe Forest across different error structures and under high-dimensional challenges.
@@
This section provides additional details for the pancreatic ductal adenocarcinoma (PDAC) re-analysis presented Section 6. We list all clinical covariates included in the analysis (Table 3) and summarize key PDAC driver genes identified in the literature (Table 4). We report the results of cross-validated model selection using the C-index in Figure 14. We show a traceplot to assess MCMC convergence for the posterior standard deviation $\sigma$ (Figure 15). Figure 16 shows the overlap in estimated propensity scores between treated and control groups.